The average running time for Broadway shows is 2 hours and 12 minutes. A producer in another city claims that the length of time of productions in his city is the same. He samples 8 shows and finds the time to be 2 hours and 5 minutes with a standard deviation of 11 minutes. Using , is the producer correct?



1
Expert's answer
2021-10-14T12:02:35-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=132\ min"

"H_1:\mu\not=132\ min"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05, df=n-1"

"=8-1=7" ​​degrees of freedom, and the critical value for a two-tailed test is"t_c =2.364619."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.364619\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{125-132}{11/\sqrt{8}}=-1.7999"

Since it is observed that "|t| = 1.7999<2.364619=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for two-tailed, "\alpha=0.05, df=7," "t=-1.7999" is "p= 0.114901," and since "p= 0.114901>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 132, at the "\alpha = 0.05" significance level.


Therefore, there is enough evidence to claim that the producer is correct, at the "\alpha = 0.05" significance level.